12. Suppose the demand curve for a product is P = 12 – 0.5Q. Which of the following is accurate?

a. The maximum willingness to pay for the 10th unit is $7.

b. If the price is $8, the quantity demanded is 4 units.

c. The maximum willingness to pay for the 8th unit is $10

. d. If the price is $6, the quantity demanded is 10 units.

e. None of the above.

Please show detail.
Answer #1

a. The maximum willingness to pay for the 10th unit is $7.

P = 12 - 0.5Q

P = 12 - 0.5(10) = 12 - 5 = $ 7

At P = 8, Q = (12 - P)/0.5 = (12 - 8)/0.5 = 4/0.5 = 8

When Q = 8 then P = 12 - 0.5(8) = 12 - 4 = 8

When P = 6 then Q = (12 - 6)/0.5 = 6/0.5 = 12